OpenStudy (calculusxy):

MEDAL! Day of the week. Price Change Monday. +0.35 Tuesday. +0.45 Wednesday. -1.56 Thursday. -0.33 What does the 0 represent in this situation? A. The price of the stock in the beginning of Monday. B. The overall price change in the week, C. The price of the stock in Friday. D. The price of the stock in the end of Thursday,

OpenStudy (calculusxy):

@hartnn

OpenStudy (calculusxy):

@amistre64

OpenStudy (amistre64):

what 0?

OpenStudy (calculusxy):

The 0 in front of the numbers.

OpenStudy (amistre64):

its a unit place value for the number of 1s

OpenStudy (calculusxy):

I want to know if it is B or not.

OpenStudy (amistre64):

the question still makes no sense

OpenStudy (calculusxy):

That's the question.

OpenStudy (amistre64):

you have no starting value to assess the price changes with ... and the only 0 there is displayed here has no bearing on anything other than how many unit values of 1 are altered during a change. for example; spose i had $5.02 at the start of the week and it changes by -0.02 5.02 -0.02 ------ 5.00 is what i have left

OpenStudy (calculusxy):

Can you help me on another one because this is not working out?

OpenStudy (calculusxy):

What is the opposite of the opposite of 8 1/2? A. +(+ 8 1/2) B. - (- 8 1/2) C. +(- 8 1/2) D. -(+ 8 1/2)

OpenStudy (amistre64):

if anything, 0 represents the initial state; and might be meant to suggest the starting value

OpenStudy (amistre64):

the opposite of A is the value x, such that A+x = 0

OpenStudy (calculusxy):

Is it C.

OpenStudy (amistre64):

no A + x = 0, when x = -A the opposite of A is -A the opposite of the opposite is .....

OpenStudy (calculusxy):

I get that it's positive, but don't get the expression.

OpenStudy (amistre64):

its not 'positive'; since we are looking at syntax lets try it this way: (-A) + x = 0 , if x = the opposite of the (opposite of A) -(-A) -(-A) ---------------- x = ______

OpenStudy (calculusxy):

I am a middle schooler. So I don't really get syntax and all that stuff.

OpenStudy (amistre64):

syntax is just the way something looks like ... how it is written as opposed to some quantitative value.

OpenStudy (amistre64):

A and B have the same quantitative value ... but differ in how they look. So when trying to determine the opposite of an object ....

OpenStudy (calculusxy):

I don't understand.

OpenStudy (calculusxy):

@amistre64

OpenStudy (anonymous):

The opposite of the opposite of something is the thing that you started with, I would think.

OpenStudy (calculusxy):

But which expression represents it?

OpenStudy (anonymous):

Well, A and B have the same result so it could be either.

OpenStudy (anonymous):

The negative times the negative equals positive 8 1/2 and the same is true for +(+8 1/2).

OpenStudy (calculusxy):

But I think is c because at first you do the opposite of eigth half and get -8 1/2 then you also do the opposite of that and get positive. So I thought maybe +(- 8 1/2)

OpenStudy (anonymous):

I think the question is quite confusing, to be perfectly honest. Does (+) equal +1 and likewise does (-) equal -1? Do they want the expression that would be an opposite times it's opposite or do they want the option that would equal the roundabout return there? Too many possibilities in the question for me to answer confidently, tbh. If this is homework, I'd ask the teacher to clarify the question for you.

OpenStudy (calculusxy):

Thank you for helpijg

OpenStudy (anonymous):

However, - 8 1/2 IS the opposite of 8 1/2. Therefore, the opposite of that opposite is the original + 8 1/2. C could possibly represent that return, as long as it's not taken in a literal mathematical sense (as a + times a negative is a negative). If you take it in a mathematical sense, then what you need the answer to be is x=8 1/2, with x being the multiple choice option. Now, both A (+(+81/2)) and B (-(+8 1/2)) equal + 8 1/2, which means either could work as an answer in that context.

OpenStudy (anonymous):

I'm sorry I can't help more, there just isn't enough contextual information for me to determine.

OpenStudy (amistre64):

lets try it this way: (-A) + x = 0 , if x = the opposite of the (opposite of A) -(-A) -(-A) ---------------- x = ______

OpenStudy (amistre64):

which option has the format -(-A) ???

OpenStudy (amistre64):

Jaeriko has just demonstrated that the value is not what they are defining as the solution; but rather the way that the solution looks.